if n=1445 rpm, calculate angular velocity in rad/sec.
1
Expert's answer
2012-06-19T09:43:24-0400
1 revolution = 2 * Pi rad
n = 1445 rpm = (1445 / 60) rps = 24.083 rps =
24.0833 * 2 * Pi rad/sec
= approx. 151.3199 rad/sec.